# Primordial Black Holes and Scalar-induced Gravitational Waves in Sneutrino Hybrid Inflation

### v2

@@ 5 Embedding in a Realistic Framework: Sneutrino Hybrid Inflation
where $\kappa$ and $(\lambda_{N})$ are dimensionless Yukawa couplings and ${M^{\prime}},M_{1}$ and $M_{\star}$ are three independent mass parameters. The superfields $\hat{N}$, $\hat{\phi}$ and $\hat{S}$ contain the bosonic components which are respectively: the singlet sneutrino $\hat{N}$ which plays the role of the inflaton; the waterfall field $\hat{\phi}$, which is not exactly at zero during inflation but slightly off due to the presence of the linear term and develops a non-zero vacuum expectation value (vev) after inflation; and the singlet field $S$ which is held at zero during and after inflation. Note that we assume a $Z_{4}$ symmetry to prevent an explicit right-handed (s)neutrino mass for the superfield containing the inflaton. Any other RH neutrinos are assumed to be singlets under the $Z_{4}$ and have large explicit masses.
The vev of the waterfall field after inflation is fixed by the first term on the right-hand side in eq. 5.1. During inflation, this term contributes a large vacuum energy to the potential. The waterfall superfield appears as $\hat{\phi}^{4}/M^{{}^{\prime}2}$ which will prevent the explicit singlet sneutrino masses via $Z_{4}$ discrete symmetry but this will be softly broken due to the presence of the linear term. The superpotential eq. 5.1 also respect the $U(1)_{R}$-symmetry and carries unit $R$-charge by assigning unit to $\hat{S}$ and $1/2$ to $\hat{N}$ The discrete subgroup of $Z_{4}$ symmetry acts as a matter parity under suitable conditions [35, 36]. The Kähler potential with non-zero F-terms during inflation is given as [34]
@@
where the dots are for higher-order terms and $m_{\text{Pl}}=2.43\times 10^{18}$ GeV is the reduced Planck mass. The field $S$ acquires a large mass and sits at zero during inflation. The F-term scalar potential is given by [34],
@@ 5.1 Comparing Model Parameters and Power Spectrum
Model $\gamma$ $\tilde{N}_{Rc}/m_{\text{Pl}}$ $\tilde{N}_{Ri}/m_{\text{Pl}}$ $\tilde{\phi}_{i}/m_{\text{Pl}}$ $N_{k}$ $n_{s}$ $r$ $T_{R}/\text{GeV}$ BP-1 $-0.017$ $0.100$ $\tilde{N}_{Rc}(1+0.001)$ $\phi_{0}$ $58$ $0.966$ $2.89\times 10^{-11}$ $5.6\times 10^{7}$ BP-2 $-0.017$ $0.095$ $\tilde{N}_{Rc}(1+0.001)$ $\phi_{0}$ $53$ $0.966$ $3.25\times 10^{-11}$ $6.3\times 10^{7}$ BP-3 $-0.017$ $0.090$ $\tilde{N}_{Rc}(1+0.001)$ $\phi_{0}$ $48$ $0.966$ $3.62\times 10^{-11}$ $6.8\times 10^{7}$
There are two major categories of these distortions: $\mu$-distortions, associated with chemical potential that occurs at early times, and Compton $y$-distortions, generated at redshifts $z\lesssim 5\times 10^{4}$. A $\mu$-distortion is associated with a Bose-Einstein distribution with $\mu\neq 0$. Currently, the COBE/FIRAS instruments put the most stringent constraints on spectral distortions, which restricts $\lvert\mu\rvert\lesssim 9.0\times 10^{-5}$ and $\lvert y\rvert\lesssim 1.5\times 10^{-5}$ at the $95\%$ C.L [39]. A PIXIE-like detector can investigate distortions with magnitudes $\mu\lesssim 2\times 10^{-8}$ and $y\lesssim 4\times 10^{-9}$ [40]. We find that our model parameters in table 2 satisfy the constraints of COBE/FIRAS. It is important to note that in the BPs table 2, although the mass scale $M$ is sub-Planckian the coupling $\beta$ is such a large number. We will see later in section 8 that it can be controlled by introducing a field transformation by the so-called $\alpha$-attractor.
@@ 6 PBH abundance
The variance, $\sigma(M_{\text{PBH}})$ of curvature perturbations ranges between $\sigma^{2}(M_{\text{PBH}})\simeq 10^{-2}-10^{-3}$ which corresponds to the critical threshold, $\delta_{c}\simeq 0.4-0.6$ [43, 44, 45, 46] to explain the entire abundance of DM. Mathematically, the variance is defined as [42],
@@
In fig. 6, the PBH abundance eq. 6.2 is demonstrated along with the different experimental constraints [47] for given parameter sets in table 2. For BP-$1$, the entire abundance of DM can be explained by PBH. The Hawking radiations may evaporate the PBHs and therefore there are some constraints in microlensing-related observations such as: CMB [48], EDGES [49], INTEGRAL [50, 51], Voyager [52], 511 keV [53], EGRB [54]; HSC (Hyper-Supreme Cam) [55], EROS [56], OGLE [57] and Icarus [58]; if PBHs accrete, there are constraints due to the CMB spectrum ref. [59]; finally the range around $M_{\odot}$ is constrained by LIGO-VIRGO-KAGRA observations on PBH-PBH merger  [60, 61, 62]). Future planned GW interferometers like Einstein Telescope (ET) and Cosmic Explorer (CE) are also expected to set limits on the PBH abundance see refs. [63, 64], these are shown in dashed lines in the plot. Future sensitivity reaches of the Nancy Roman Telescope from micro-lensing are also presented, see ref. [65].
@@ 7 Scalar-induced GWs
Several astrophysical sources of GW background are expected while searching for stochastic GW of cosmic origin. They can mainly be in the form of LIGO/VIRGO observed binary black hole (BH-BH) merging events [77, 78] as well as binary neutron star (NS-NS) events [79]. The foreground NS and BH may be subtracted with the sensitivities of the BBO and ET / CE windows, particularly in the range $\Omega_{\rm GW}\simeq 10^{-15}$ [80] and $\Omega_{\rm GW}\simeq 10^{-13}$ [81] to distinguish with the scalar-induced GWs of cosmic origin. The binary white dwarf galactic and extra-galactic (WD-WD) may be more significant than the NS-NS and BH-BH foregrounds in the LISA window [82] and can be subtracted [83] with the expected sensitivity at $\Omega_{\rm GW}\simeq 10^{-13}$ [84]. In addition to this subtraction, the GW spectrum produced by the astrophysical foreground increased with frequency $\propto f^{2/3}$ [85]. This GW spectrum differs from that produced by second-order gravitational waves, which at low frequencies convict as $f^{3/2}$ and at higher frequencies as $f^{-3/2}$. This will allow us to identify the GW signals generated by scalar-induced sources.
@@ 8 Model22-2- 2: $\alpha$-attractor sneutrino hybrid inflation
Along the valley, as long as $\tilde{\upsilon}_{R}\lesssim\tilde{\upsilon}_{Rc}$, the effective mass square of $\tilde{\phi}$ becomes negative. This gives rise to tachyonic instability that will grow the curvature perturbations which further enhance the scalar power spectrum at small scales. The collapse of large density fluctuations upon horizon reentry produces the PBHs. The presence of the linear term in the potential eq. 8.2 will not keep the field $\tilde{\phi}$ exactly at $\tilde{\phi}=0$ but is displaced depending upon the sign. of the coefficient of the linear term $d$. This inflates unnecessary topological defects and controls the peak of the power spectrum at small scales to avoid PBH overproduction. We took $M\sim O(1)$ to avoid eternal inflation [21]. The parameter $\gamma$ controls the amplitude of the plateau in the valley, the coupling $\lambda_{N}$ defines the number of e-folds in the waterfall regime and $\kappa$ will fix the amplitude of the power spectrum at the pivot scale $k_{\star}=0.05\,\text{Mpc}^{-1}$ that is $A_{s}\simeq 2.24\times 10^{-9}$. We consider the exponential $\alpha$-attractor model [89] to avoid the problem of initial conditions that we usually have in a standard hybrid scenario [19] (see [21, 13] for a detailed discussion of initial conditions). Considering all these constraints, we define the BPs in table 3 for the potential in eq. 8.2.
@@
Model $\gamma$ $\sqrt{\alpha}/m_{\text{Pl}}$ $\tilde{\upsilon}_{Ri}/m_{\text{Pl}}$ $\tilde{\phi}_{i}/m_{\text{Pl}}$ $N_{k}$ $n_{s}$ $r$ $T_{R}/\text{GeV}$ BP-1 $0.073$ $1$ $3.5$ 00 $60$ $0.961$ $0.0068$ $9.8\times 10^{8}$ BP-2 $0.085$ $1$ $4.0$ 00 $62$ $0.969$ $0.0056$ $9.2\times 10^{8}$
@@ 9 Fine-tuning Estimate
In this framework, $x\in\{M_{1},\beta,\gamma,\lambda_{N}\}$ in eq. 9.1. Evaluating numerically, fine-tuning estimates for sneutrino theory parameters eq. 8.2 are given in table 5. The maximum fine-tuning we obtain is $8$, which is almost five orders of magnitude smaller than single-field inflation [90] and one order of magnitude smaller than standard hybrid inflation [14].
@@
$\Delta_{x}$ $\Delta_{M_{1}}$ $\Delta_{\beta}$ $\Delta_{\gamma}$ $\Delta_{\lambda_{N}}$ FT $8$ $3$ $1$ $4$
@@ 10 Discussion and Conclusion
  • Second-order tensor perturbations propagating as GWs are predicted with amplitude $\Omega_{\rm GW}h^{2}$ $\sim 10^{-9}$ and peak frequency $f\sim 0.1$ Hz by LISA and $\Omega_{\rm GW}h^{2}\sim 10^{-11}$ and peak frequency of $\sim 10$ Hz in ET (see fig. 7). Production of PBH of mass around $10^{-13}M_{\odot}$ as the sole DM candidate in the universe is proposed. This novel DM candidate is also a signature of the sneutrino (see fig. 11).
  • The fine-tuning for this model is almost $O(1)$ or smaller (see table 4) which shows its supremacy over the single field inflationary model.
@@
  • We achieved acceptable values for both spectral index ($n_{s}\simeq 0.961-0.969$) and tensor-to-scalar ratio ($r\simeq 0.0056-0.0068$) satisfying PBH as dark matter and detectable GW signal.
  • Second-order tensor perturbations propagating as GWs are predicted with amplitude $\Omega_{\rm GW}h^{2}$ $\sim 10^{-9}$ and peak frequency f $\sim 0.1$ Hz by LISA and $\Omega_{\rm GW}h^{2}\sim 10^{-11}$ and peak frequency of $\sim 10$ Hz in ET in this model (see fig. 12). Production of PBH of mass around $10^{-13}M_{\odot}$ as the sole DM candidate in the universe is proposed. This novel DM candidate is also a signature of the sneutrino (see fig. 11).
@@
Thus, we offer one potential way to test the origin of right-handed neutrino mass generation, which is currently inaccessible in colliders, using a GW detector.
As a future outlook, it could be interesting to study the impact of non-Gaussianities during the waterfall transition in the models studied as they impact the abundance of PBH formation rate, PBH clustering, and the amplitude of scalar-induced GW signals (see ref. [91, 92] for recent reviews). If some characteristic features of the GW spectral shapes encountered in this study are observed, one may look to target additional observations to distinguish between SUSY-mediated inflation and other scenarios. Particularly in sneutrino masses of TeV, these could be searched in experiments [93] at the particle physics laboratories. In this manner, we can complement GW searches with laboratory searches in the same BSM parameter space. Gravitational wave astronomy with the planned global network of GW detectors aspires to achieve measurement precisions that are orders of magnitude better than the present day detectors. This new era of GW detectors worldwide will make the dream of testing fundamental BSM mechanisms, e.g. for supersymmetry physics, or neutrino physics of the universe and inflationary cosmology, a reality in the near future.

### v4

@@ 5 Embedding in a Realistic Framework: Sneutrino Hybrid Inflation
where $\kappa$ and $(\lambda_{N})$ are dimensionless Yukawa couplings and ${M^{\prime}},M_{1}$ and $M_{\star}$ are three independent mass parameters. The superfields $\hat{N}$, $\hat{\phi}$ and $\hat{S}$ contain the bosonic components which are respectively: the singlet sneutrino $\hat{N}$ which plays the role of the inflaton; the waterfall field $\hat{\phi}$, which is not exactly at zero during inflation but slightly off due to the presence of the linear term and develops a non-zero vacuum expectation value (vev) after inflation; and the singlet field $S$ is held at zero during and after inflation. Note that we assume a $Z_{4}$ symmetry to prevent an explicit right-handed (s)neutrino mass for the superfield containing the inflaton. Any other RH neutrinos are assumed to be singlets under the $Z_{4}$ and have large explicit masses.
The vev of the waterfall field after inflation is fixed by the first term on the right-hand side in eq. 5.1. During inflation, this term contributes a large vacuum energy to the potential. The waterfall superfield appears as $\hat{\phi}^{4}/M^{{}^{\prime}2}$ which will prevent the explicit singlet sneutrino masses via $Z_{4}$ discrete symmetry but this will be softly broken due to the presence of the linear term. The superpotential eq. 5.1 also respect the $U(1)_{R}$-symmetry and carries unit $R$-charge by assigning unit to $\hat{S}$ and $1/2$ to $\hat{N}$. The discrete subgroup of $Z_{4}$ symmetry acts as a matter parity under suitable conditions [56, 57]. The Kähler potential with non-zero F-terms during inflation is given as [55]
@@
where the dots are for higher-order terms  [^4] Equation 2.1 is just used as a toy model without committing to a particle theory scenario like SUSY, etc. The results obtained in this toy model help us study the (s)neutrino model presented in eq. 5.5 based on SUSY. For instance, the terms containing dimension-6 in eq. 2.1 in the realistic SUSY model are allowed under the assumed $Z_{4}$-symmetry. Such terms in SUSY (and SUGRA) with the cut-off scale $M_{\star}$ are perfectly consistent as the model is assumed to be valid up to this energy scale and not beyond that. For more details of such discussion see the original model presented in [55]. and $m_{\text{Pl}}=2.43\times 10^{18}$ GeV is the reduced Planck mass. The field $S$ acquires a large mass and sits at zero during inflation. The F-term scalar potential is given by [55],
@@ 5.1 Comparing Model Parameters and Power Spectrum
Model $\gamma$ $\tilde{N}_{Rc}/m_{\text{Pl}}$ $\tilde{N}_{Ri}/m_{\text{Pl}}$ $\tilde{\phi}_{i}/m_{\text{Pl}}$ $N_{k}$ $n_{s}$ $r$ $T_{R}/\text{GeV}$ $\delta_{c}$ BP-1 $-0.017$ $0.100$ $\tilde{N}_{Rc}(1+0.001)$ $\phi_{0}$ $58$ $0.966$ $2.89\times 10^{-11}$ $5.6\times 10^{7}$ $0.52$ BP-2 $-0.017$ $0.095$ $\tilde{N}_{Rc}(1+0.001)$ $\phi_{0}$ $53$ $0.966$ $3.25\times 10^{-11}$ $6.3\times 10^{7}$ $0.45$ BP-3 $-0.017$ $0.090$ $\tilde{N}_{Rc}(1+0.001)$ $\phi_{0}$ $48$ $0.966$ $3.62\times 10^{-11}$ $6.8\times 10^{7}$ $0.6$
There are two major categories of these distortions: $\mu$-distortions, associated with chemical potential that occurs at early times, and Compton $y$-distortions, generated at redshifts $z\lesssim 5\times 10^{4}$. A $\mu$-distortion is associated with a Bose-Einstein distribution with $\mu\neq 0$. Currently, the COBE/FIRAS instruments put the most stringent constraints on spectral distortions, which restricts $\lvert\mu\rvert\lesssim 9.0\times 10^{-5}$ and $\lvert y\rvert\lesssim 1.5\times 10^{-5}$ at the $95\%$ C.L [60, 61, 62]. A PIXIE-like detector can investigate distortions with magnitudes $\mu\lesssim 2\times 10^{-8}$ and $y\lesssim 4\times 10^{-9}$ [63]. The power spectrum on small scales, should be around $10^{-2}-10^{-3}$ to explain the maximum abundance of DM. While the peak is constrained from FIRAS on large scales see the bounds in fig. 5. We find that our model parameters in table 2 satisfy the constraints of COBE/FIRAS. It is important to note that in the BPs table 2, although the mass scale $M$ is sub-Planckian the coupling $\beta$ is such a large number. We will see later in section 8 that it can be controlled by introducing a field transformation by the so-called $\alpha$-attractor.
@@ 6 PBH abundance
The variance, $\sigma(M_{\text{PBH}})$ of curvature perturbations ranges between $\sigma^{2}(M_{\text{PBH}})\simeq 10^{-2}-10^{-3}$ which corresponds to the critical threshold, $\delta_{c}\simeq 0.4-0.6$ [66, 67, 68, 69] to explain the entire abundance of DM. Although $\delta_{c}$ is not a free parameter, it depends on both the shape of the power spectrum [66] and the choice of the window function [70], still there remains some uncertainty in the literature regarding the calculation of the threshold, linked to super-horizon corrections [71] and the average profile for the compaction function [72]. In our analysis, we neglect the non-linear relationship between the density contrast and the curvature perturbation field and approximate the threshold without accounting for its dependence on the shape of the power spectrum and the window function; rather we have chosen it in such a way that it can explain the PBH as DM up to the allowed fraction in a given mass range. The corresponding $\delta_{c}$ values for each BPs are given in table 2 and table 3. Note that the parameter space is expected to change in a non-trivial way when moving beyond these approximations, particularly concerning corrections arising from non-linearities.  [^5] Note that there exists a non-linear relation between curvature perturbation and density contrast field that may have some impact on the PBH abundance which can be circumvented by the uncertainties in parameters like $\gamma_{c}$. We have discussed this in detail in the Appendix appendix A. Mathematically, the variance is defined as [65],
@@
In fig. 6, the PBH abundance eq. 6.2 is demonstrated along with the different experimental constraints [73] for given parameter sets in table 2. Note that we consider the case of monochromatic mass functions. The bounds may change in fig. 6 and fig. 11 for an extended mass function [74]. For BP-$1$, the entire abundance of DM can be explained by PBH. The Hawking radiations may evaporate the PBHs and therefore there are some constraints in microlensing-related observations such as: CMB [75], EDGES [76], INTEGRAL [77, 78], Voyager [79], 511 keV [80], EGRB [81]; HSC (Hyper-Supreme Cam) [82], EROS [83], OGLE-IV [84, 85] and Icarus [86]; if PBHs accrete, there are constraints due to the CMB spectrum ref. [87]; finally the range around $M_{\odot}$ is constrained by LIGO-VIRGO-KAGRA observations on PBH-PBH merger  [88, 89, 90]). Future planned GW interferometers like LISA, Einstein Telescope (ET) and Cosmic Explorer (CE) are also expected to set limits on the PBH abundance see refs. [91, 92, 93], these are shown in dashed lines in the plot. Future sensitivity reaches of the Nancy Roman Telescope from micro-lensing are also presented, see ref. [94] [^6] A recent study of the sensitivity of X-ray telescopes studying x-ray pulsars to microlensing by sub-atomic size primordial black holes is done in [95]..
@@ 7 Scalar-induced GWs
Searching for stochastic GW of cosmic origin is likely to reveal multiple astrophysical sources of GW background. They can primarily take the form of binary neutron star (NS-NS) events [107] and LIGO/VIRGO detected binary black hole (BH-BH) merging events [108, 109]. To differentiate with the scalar-induced GWs of cosmic origin, the foreground NS and BH can be subtracted using the sensitivities of the BBO and ET / CE windows, especially in the range $\Omega_{\rm GW}\simeq 10^{-15}$ [110] and $\Omega_{\rm GW}\simeq 10^{-13}$ [111]. The binary white dwarf galactic and extra-galactic (WD-WD) can be removed [112] with the expected sensitivity at $\Omega_{\rm GW}\simeq 10^{-13}$ [113], and may be more significant than the NS-NS and BH-BH foregrounds in the LISA window [114]. The GW spectrum generated by the astrophysical foreground grew with frequency $\propto f^{2/3}$ in addition to this subtraction [115]. This GW spectrum differs from that produced by second-order gravitational waves, which at low frequencies convict as $f^{3/2}$ and at higher frequencies as $f^{-3/2}$. This will allow us to identify the GW signals generated by scalar-induced sources.
@@ 8 Model22-2- 2: $\alpha$-attractor sneutrino hybrid inflation
First we discuss what happens along the valley: as long as $\tilde{\upsilon}_{R}\lesssim\tilde{\upsilon}_{Rc}$, the effective mass square of $\tilde{\phi}$ stays negative. This gives rise to kind of tachyonic instability that leads to the large growth the curvature perturbations which consequently further enhance the scalar power spectrum at small scales of the universe. The collapse of large density fluctuations upon horizon reentry leads to PBH production. The presence of the linear term in the potential eq. 8.2 serves as not to keep the field $\tilde{\phi}$ exactly at $\tilde{\phi}=0$ but makes it displaced. This also depends upon the sign of the coefficient of the linear term $d$. On the one hand, this inflates away unnecessary topological defects and on the other hand also controls the peak of the power spectrum at small scales to rescues from PBH overproduction. We take $M\sim O(1)$ to avoid eternal inflation [23]. The parameter $\gamma$ controls the amplitude of the plateau in the valley, the coupling $\lambda_{N}$ defines the number of e-folds in the waterfall regime and $\kappa$ will fix the amplitude of the power spectrum at the pivot scale $k_{\star}=0.05\,\text{Mpc}^{-1}$ that is $A_{s}\simeq 2.24\times 10^{-9}$. We consider the model involving exponential $\alpha$-attractor [119] in order to to get rid of the problem of initial conditions that we usually have in a standard hybrid scenario [21] (see [23, 13] for a detailed discussion of such initial conditions). Having considered these constraints, we define the BPs in table 3 for the potential in eq. 8.2.
@@
Model $\gamma$ $\sqrt{\alpha}/m_{\text{Pl}}$ $\tilde{\upsilon}_{Ri}/m_{\text{Pl}}$ $\tilde{\phi}_{i}/m_{\text{Pl}}$ $N_{k}$ $n_{s}$ $r$ $T_{R}/\text{GeV}$ $\delta_{c}$ BP-1 $0.073$ $1$ $3.5$ 00 $60$ $0.961$ $0.0068$ $9.8\times 10^{8}$ $0.524$ BP-2 $0.085$ $1$ $4.0$ 00 $62$ $0.969$ $0.0056$ $9.2\times 10^{8}$ $0.460$
@@ 9 Fine-tuning Estimate
In this framework, $x\in\{M,\beta,\gamma,\lambda_{N}\}$ in eq. 9.1. Evaluating numerically, fine-tuning estimates for sneutrino theory parameters eq. 8.2 are given in table 5. The maximum fine-tuning we obtain is $8$, which is almost five orders of magnitude smaller than single-field inflation [120] and one order of magnitude smaller than standard hybrid inflation [14].
@@
$\Delta_{x}$ $\Delta_{M}$ $\Delta_{\beta}$ $\Delta_{\gamma}$ $\Delta_{\lambda_{N}}$ FT $8$ $3$ $1$ $4$
@@ 10 Discussion and Conclusion
  • Second-order tensor perturbations propagating as GWs are predicted with amplitude $\Omega_{\rm GW}h^{2}$ $\sim 10^{-9}$ and peak frequency $f\sim 0.1$ Hz by LISA and $\Omega_{\rm GW}h^{2}\sim 10^{-11}$ and peak frequency of $\sim 10$ Hz in ET (see fig. 7). The production of PBH of mass around $10^{-13}M_{\odot}$ as the sole DM candidate in the universe is proposed. This novel candidate for DM is also a signature of the sneutrino (see fig. 11).
  • The fine-tuning for this model is almost $O(1)$ or smaller (see table 4) which shows its supremacy over the single-field inflationary model.
@@
  • We achieved acceptable values for both the spectral index ($n_{s}\simeq 0.961-0.969$) and tensor-to-scalar ratio ($r\simeq 0.0056-0.0068$) satisfying PBH as dark matter and detectable GW signal.
  • Second-order tensor perturbations propagating as GWs are predicted with amplitude $\Omega_{\rm GW}h^{2}$ $\sim 10^{-9}$ and peak frequency f $\sim 0.1$ Hz by LISA and $\Omega_{\rm GW}h^{2}\sim 10^{-11}$ and peak frequency of $\sim 10$ Hz in ET in this model (see fig. 12). The production of PBH of mass around $10^{-13}M_{\odot}$ as the sole DM candidate in the universe is proposed. This novel candidate for DM is also a signature of the sneutrino (see fig. 11).
@@
Thus, we offer one potential way to test the origin of right-handed neutrino mass generation, which is currently inaccessible in colliders, using a GW detector. Thus we find that if the cosmic inflation is driven by the sneutrino responsible generation of right-handed neutrino mass, one could see its imprints in gravitational lensing observations and PBH as the entire DM candidate and be tested in upcoming GW experiments.
As a future outlook, it could be interesting to study the impact of non-Gaussianities during the waterfall transition in the models studied as they impact the abundance of PBH formation rate, PBH clustering, and the amplitude of scalar-induced GW signals (see ref. [121, 122] for recent reviews). If some characteristic features of the GW spectral shapes encountered in this study are observed, one may look to target additional observations to distinguish between SUSY-mediated inflation and other scenarios. Particularly in sneutrino masses of TeV, these could be searched in experiments [123] at the particle physics laboratories. In this manner, we can complement GW searches with laboratory searches in the same BSM parameter space. Gravitational wave astronomy with the planned global network of GW detectors aspires to achieve measurement precisions that are orders of magnitude better than the present-day detectors. This new era of GW detectors worldwide will make the dream of testing fundamental BSM mechanisms, e.g. for supersymmetry physics, or neutrino physics of the universe and inflationary cosmology, a reality in the near future.